OpenStudy (anonymous):

How do I solve this? tan(theta - 10) = cot(2 theta + 17)

OpenStudy (khaakenajaf):

cos(A-B) = cos(A)cos(B) + sin(A)sin(B) sin(A-B) = sin(A)cos(B) -cos(A)sin(B) divide them A=theta B=10 i hope it will help

OpenStudy (akashdeepdeb):

^ That maybe effective, there is also a method which goes like this. :) cot A = tan (90 - A) \[cot(2\theta+17) = tan(90 - (2\theta+17))\] And this you know is now \( = tan(\theta - 10)\) Thus \( = tan(\theta - 10)\) = \(tan(90 - (2\theta+17))\) NOTE: If tan A = tan B , this implies that, A=B Now can you find \(\theta\) ? :)

OpenStudy (anonymous):

tan (x - 10) = cot (- 2x + 17) = tan (90 - 2x - 17) = tan (-2x + 63) a) (x - 10) = -2x + 63 -> 3x = 73 -> x = 73/3 degree = 24.20'' b) Add another answer: tan x = tan (x + 180) tan (x - 10) = tan (-2x + 63 + 180) --> x - 10 = - 2x + 243 --> 3x = 253 --> x = 253/3 degree